# Higgs-Confinement Transitions

### v2

@@ 1.3 Application to QCD at Finite Baryon Density
As our final example, we consider conventional three-flavor QCD (without elementary Higgs fields), with real, positive quark masses. (See section 6 for a full account.) We often simplify the discussion by assuming a common positive quark mass $m_{q}>0$, though some of our results are more general and also apply for physical quark masses. QCD with three Dirac flavors preserves ${\mathsf{C}}$, ${\mathsf{P}}$, and ${\mathsf{T}}$, as well as $U(1)_{B}$ baryon number and (in the case of a common quark mass) a vector-like $SU(3)_{f}$ flavor symmetry.
@@ 2.1 What is an SPT Phase?
In this section we review the notion of Symmetry Protected Topological (SPT) phases, elaborating on section 1.2.1 (which contains many references). A typical SPT is a gapped theory that is essentially trivial: it has a unique vacuum, and no dynamical degrees of freedom – massless or topological – at long distances. In condensed matter parlance it has neither conventional nor topological order.484848 In particular, it has a one-dimensional Hilbert space on any spatial manifold. This implies that all global symmetries – continuous or discrete, ordinary or generalized – are unbroken. In the Landau paradigm, such a phase is considered to be trivial.
@@
  • In this paper we will restrict to spacetime manifolds that are orientable and carry a spin structure. The theories we discuss can also be analyzed on manifolds that do not admit such structures,494949 For unorientable manifolds this requires a choice of orientation-reversing symmetry such as time-reversal or parity; for manifolds that do not admit a spin structure, it may require a twist by a suitable internal symmetry, such as baryon number. SPTs for the flavor and time-reversal symmetries of QCD on unorientable manifolds were classified in [102]. which leads to stronger results that we will present in [70].
@@
The partition function $Z[g,A]$, obtained by path integrating over all dynamical fields, is a functional of the background fields.505050 It also depends on the topology of $M_{4}$, as well as the orientation and the spin structure. Because the theory is gapped, it is a local functional, which can be written in terms of a local effective action $S_{\text{eff}}[g,A]$ for the background fields. In Euclidean signature,
@@
    Topological terms in the background fields, with quantized coefficients. These are typically purely imaginary in Euclidean signature, i.e. they only affect the phase of the partition function.515151 If one chooses to only keep the imaginary topological terms in $S_{\text{eff}}$, the partition function is a pure phase, whose complex conjugate coincides with its inverse. The resulting theories are known as invertible, and their relation to SPTs is discussed in [53]. They can still be thought of as well-defined local counterterms, albeit rigid ones that cannot be deformed continuously.
@@
    We will refer to all such quantized anomaly-inflow terms as SPTs, though this use is somewhat broader then in the condensed matter literature.525252 The term SPT is typically reserved for situations that require a particular global symmetry. Some invertible anomaly-inflow actions do not require any such symmetry, see for instance [103, 104, 51, 105, 106, 107]. The preceding discussion makes it clear that SPTs can be usefully approached from two complementary points of view: via the anomaly-inflow terms for background fields in the bulk, or in terms of the anomalous boundary degrees of freedom. In this paper we mostly take the former, bulk point of view.
@@
For our purposes in this paper, the most important thing about the Chern-Simons SPT in (2.2) is the quantization of its coefficient $k\in{\mathbb{Z}}$. If we imagine dialing the continuous coupling constants of the theory while preserving the $U(1)$ symmetry, and without closing the bulk gap,545454 Recall that a boundary, if present, necessarily hosts gapless degrees of freedom to cancel the bulk anomaly. then $k$ does not change. There are only two ways in which $k$ can change:
@@ 2.2.1 Basic Properties
  • The $\theta_{f}=\pi$ SPT represents the anomaly-inflow action for the parity anomaly [108, 109] in 2+1 dimensions: on a manifold $M_{4}$ with boundary $M_{3}=\partial M_{4}$, the $\theta_{f}=\pi$ SPT can be integrated by parts to produce a Chern-Simons term (2.2) with level $k={1\over 2}$ on $M_{3}$.555555 This in turn gives rise to a half-integer Hall conductance in units of $e^{2}/h$, as discussed in section 2.2.3 below. Since this term flips sign under either ${\mathsf{T}}$ or $\sf C{\mathsf{T}}$, the partition function of the boundary degrees of freedom on $M_{3}$ must compensate by transforming as follows,
@@ 2.2.2 Example: Massive Dirac Fermion
Here $\alpha=1,2$ is a left-handed Weyl spinor index.565656 We use the conventions of Wess and Bagger [65] for 2-component Weyl spinors. See also [74]. Hermitian conjugation exchanges left-handed undotted Weyl spinors with dotted right-handed Weyl spinors. Thus the hermitian conjugate fields are
@@
    Thanks to the standard axial-vector triangle anomaly, the $U(1)_{\text{axial}}$ symmetry is anomalous if we turn on a background field $A$ for the vector-like $U(1)$ flavor symmetry.585858 For now we ignore gravitational backgrounds; we will include them below. This leads to the following transformation rule for the partition function,595959 This formula can be derived using the standard descent formalism, which relates triangle anomalies to index theorems; see for instance section 2 of [11] for a recent pedagogical discussion with references.
@@ 2.2.3 Detecting the $\theta_{f}=\pi$ SPT via 2+1d Quantum Hall Conductance
Consider the codimension-one defect that generates the ${\mathsf{T}}$-symmetry, oriented along a 2+1d submanifold of spacetime. Since this defect implements the action of ${\mathsf{T}}$, it maps the $\theta_{f}=\pi$ SPT on one side to $\theta_{f}=-\pi$, leading to an overall jump by $\Delta\theta_{f}=2\pi$ across the wall. This has the effect of decorating the ${\mathsf{T}}$-defect by a $U(1)$ Chern-Simons term with level $k=1$,606060 See [119, 52, 120, 121, 122, 123] for a discussion of decorated domain walls. which in turn implies a non-zero quantum Hall conductance $\sigma_{xy}=1$, in units of $e^{2}/h$ (with $e$ the electron charge, and $h$ Planck’s constant), see for instance [39, 124].
@@ 2.3.1 Basic Properties
On a manifold with boundary this SPT can be integrated by parts to produce what is effectively a level-${1\over 2}$ gravitational Chern-Simons term,616161 This in turn gives rise to a thermal Hall conductance $\kappa_{xy}=1/4+n/2$ with $n\in{\mathbb{Z}}$, in units of $\pi^{2}k_{B}^{2}T/3h$, where $T$ is temperature, $h$ is Planck’s constant, and $k_{B}$ is Boltzmann’s constant. This is further discussed in section 2.3.3 below. whose flip under ${\mathsf{T}}$ or ${\mathsf{C}}{\mathsf{T}}$ transforms the partition function by a properly quantized gravitational Chern-Simons counterterm. This is the purely gravitational version of the parity anomaly.
@@ 2.3.2 Example: Massive Weyl Fermion
Note that a single Weyl fermion does not have charge-conjugation symmetry.626262 The ${{\mathsf{C}}{\mathsf{P}}{\mathsf{T}}}$ theorem also guarantees invariance under ${\mathsf{C}}{\mathsf{P}}:\psi_{\alpha}(t,\vec{x})\to i\overline{\psi}^{\dot{\alpha}}(t,-\vec{x})$. As before, we choose counterterms so that $m>0$ corresponds to the trivial SPT.
@@ 2.5.2 Partition Functions and Gapless SPTs
One way to diagnose gapped SPT phases is via their partition function in the presence of background fields on euclidean spacetime manifolds. Here we will make some related observations in the context of gapless SPTs.636363 We are grateful to Ryan Thorngren for discussions about the content of this subsection, and for explaining some of the intuitive arguments underlying [87]. We distinguish three cases:
@@ 2.5.3 Gapless SPT Phase with a $U(1)$ Nambu-Goldstone Boson
In this paper we will encounter gapless phases that contain a NGB for a spontaneously broken flavor symmetry $U(1)_{B}$ with background field $B$. In addition, there can be SPTs for a distinct, unbroken $U(1)$ flavor symmetry with background field $A$, as well as gravity, both protected by an unbroken time-reversal symmetry. Let us take the $U(1)_{B}$ Goldstone boson $\phi$ to be a dimensionless scalar of periodicity $\phi\sim\phi+2\pi$. The shift symmetry of $\phi$, with current $f_{\phi}d\phi$ coupling to $B$, is the broken $U(1)_{B}$ symmetry.646464 Here $f_{\phi}$ is the decay constant of $\phi$. In addition, $\phi$ has an unbroken, generally emergent, $U(1)_{C}^{(2)}$ two-form winding symmetry, whose current $\sim*d\phi$ couples to a three-form gauge field $C^{(3)}$ with field strength $H^{(4)}=dC^{(3)}$. All together the low-energy Lagrangian for $\phi$ is,
@@
Note that the broken $U(1)_{B}$ symmetry and the unbroken $U(1)_{C}^{(2)}$ winding symmetry have a mixed ’t Hooft anomaly. In the presentation (2.53), the theory is manifestly invariant under $C^{(3)}$ background gauge transformations,656565 If we add to the action a counterterm $\sim\int_{M_{4}}B\wedge C^{(3)}$ then a $C^{(3)}$ gauge transformation leads to a shift $\sim dB$. Clearly such a shift cannot eat an SPT involving only the $U(1)$ background field $A$ or gravity. but a $U(1)_{B}$ background gauge transformation $B\to B+d\lambda_{B}$ shifts
@@
Note that both ${1\over 2\pi}H^{(4)}$ and ${1\over 8\pi^{2}}F\wedge F$ have integral periods. However, substituting back into (2.54) we find a non-vanishing mixed anomaly between the broken $U(1)_{B}$ symmetry and the unbroken $U(1)$, in accord with the general discussion in 2.5.1. Completely analogously, comparing with (2.27), we find that we can absorb $\theta_{g}=\pi$ SPT by choosing
@@ 3.1 Fields and Lagrangian
Here $\alpha=1,2$ is a left-handed Weyl spinor index,666666 We use the conventions of Wess and Bagger [65] for 2-component Weyl spinors. See [74] for a presentation of QCD in these conventions. which we will generally suppress. The raised and lowered $a$-indices are fundamental and anti-fundamental $SU(N)_{c}$ color indices, respectively, so that the pair $(\psi^{a}_{i},\chi_{a}^{i})$ transforms as $({\bf N},\overline{\bf N})$ of the $SU(N)_{c}$ gauge group. Likewise, the lowered and raised $i$-indices are fundamental and anti-fundamental $SU(N_{f})$ flavor indices, and $(\psi^{a}_{i},\chi_{a}^{i})$ transforms as $({\bf N_{f}},\overline{\bf N_{f}})$ under $SU(N_{f})$.
@@
In this paper we will mostly discuss the case of $N_{f}$ degenerate quarks, with a common real Dirac mass $m_{q}\in{\mathbb{R}}$, and vanishing theta-angle, $\theta_{c}=0$, for the dynamical $SU(N)_{c}$ gauge field.676767 An anomalous $U(1)_{\text{axial}}$ transformation can be used to eliminate $\theta_{c}$ in favor of a (generally) complex mass $m_{q}\in{\mathbb{C}}$. Below we will additionally impose time-reversal symmetry, which requires real $m_{q}\in{\mathbb{R}}$. The Lagrangian in Lorentzian signature is then given by
@@ 3.2 Continuous Symmetries
The continuous gauge and global symmetries that act faithfully on the quark fields are686868 See for instance section 4.1 of [113] for a related discussion.
@@
Here $SU(N)_{c}$ is the gauge group and $U(N_{f})_{f}$ is the vector-like flavor symmetry preserved by the quark mass.696969 Whenever we wish to highlight that something pertains to color or flavor we will use the subscripts $c$ and $f$ respectively, though we will omit them if confusion is unlikely to arise. Note that
@@
    Such a theory is bosonic: even though it is formulated using fermionic quark fields, all gauge invariant local operators are bosons and the theory can be studied on non-spin manifolds.707070 This utilizes the symmetry group $(SU(N)_{c}\times\text{Spin}(4)_{\text{Lorentz}})/{\mathbb{Z}}_{2}^{F}$ that arises when $N$ is even. On oriented Riemannian manifolds $M_{4}$ without a spin structure, the second Stiefel-Whitney class $w_{2}(M_{4})\in H^{2}(M_{4},{\mathbb{Z}}_{2})$ of the tangent bundle is non-trivial, while the gauge fields are connections on an $SU(N)_{c}/{\mathbb{Z}}_{2}$ bundle. The ${\mathbb{Z}}_{2}^{F}$ quotient in the symmetry group implies that the class that measures the obstruction to lifting the $SU(N)_{c}/{\mathbb{Z}}_{2}$ gauge bundle to an $SU(N)_{c}$ bundle is precisely given by $w_{2}(M_{4})$.
@@
    In particular this implies that $(-1)^{F}$ is identified with $(-1)\in U(1)_{B}\subset G_{f}$, so that fundamental baryons and anti-baryons with $B=\pm 1$ are necessarily fermions. This is consistent with the fact that they contain an odd number $N$ of constituent quarks. This identification further implies that the theory can be placed on non-spin manifolds, as long as the $U(1)_{B}$ background gauge field is taken to be a spinc connection.717171 In general this leads to additional information not accessible on spin manifolds, see e.g. [110, 139, 140, 141, 142].

### v3

@@ 1.3 Application to QCD at Finite Baryon Density
As our final example, we consider conventional three-flavor QCD (without elementary Higgs fields), with real, positive quark masses. (See section 6 for a full account.) We often simplify the discussion by assuming a common positive quark mass $m_{q}>0$, though some of our results are more general and also apply for physical quark masses. QCD with three Dirac flavors preserves ${\mathsf{C}}$, ${\mathsf{P}}$, and ${\mathsf{T}}$, as well as $U(1)_{B}$ baryon number and (in the case of a common quark mass) a vector-like $SU(3)_{f}$ flavor symmetry.484848 Famously, QCD admits a $\theta$-angle for its $SU(3)$ gauge group. (Note that the value of $\theta$ modulo $2\pi$ is meaningful, since we have fixed the quark masses to be positive.) However, experiments constrain $|\theta|\lesssim 10^{-10}$ (see chapter 9 of the PDG [93]). Thus taking $\theta\simeq 0$, so that ${\mathsf{T}}$-symmetry is neither explicitly nor spontaneously broken, is an excellent approximation.
@@ 2.1 What is an SPT Phase?
In this section we review the notion of Symmetry Protected Topological (SPT) phases, elaborating on section 1.2.1 (which contains many references). A typical SPT is a gapped theory that is essentially trivial: it has a unique vacuum, and no dynamical degrees of freedom – massless or topological – at long distances. In condensed matter parlance it has neither conventional nor topological order.494949 In particular, it has a one-dimensional Hilbert space on any spatial manifold. This implies that all global symmetries – continuous or discrete, ordinary or generalized – are unbroken. In the Landau paradigm, such a phase is considered to be trivial.
@@
  • In this paper we will restrict to spacetime manifolds that are orientable and carry a spin structure. The theories we discuss can also be analyzed on manifolds that do not admit such structures,505050 For unorientable manifolds this requires a choice of orientation-reversing symmetry such as time-reversal or parity; for manifolds that do not admit a spin structure, it may require a twist by a suitable internal symmetry, such as baryon number. SPTs for the flavor and time-reversal symmetries of QCD on unorientable manifolds were classified in [103]. which leads to stronger results that we will present in [70].
@@
The partition function $Z[g,A]$, obtained by path integrating over all dynamical fields, is a functional of the background fields.515151 It also depends on the topology of $M_{4}$, as well as the orientation and the spin structure. Because the theory is gapped, it is a local functional, which can be written in terms of a local effective action $S_{\text{eff}}[g,A]$ for the background fields. In Euclidean signature,
@@
    Topological terms in the background fields, with quantized coefficients. These are typically purely imaginary in Euclidean signature, i.e. they only affect the phase of the partition function.525252 If one chooses to only keep the imaginary topological terms in $S_{\text{eff}}$, the partition function is a pure phase, whose complex conjugate coincides with its inverse. The resulting theories are known as invertible, and their relation to SPTs is discussed in [53]. They can still be thought of as well-defined local counterterms, albeit rigid ones that cannot be deformed continuously.
@@
    We will refer to all such quantized anomaly-inflow terms as SPTs, though this use is somewhat broader then in the condensed matter literature.535353 The term SPT is typically reserved for situations that require a particular global symmetry. Some invertible anomaly-inflow actions do not require any such symmetry, see for instance [104, 105, 51, 106, 107, 108]. The preceding discussion makes it clear that SPTs can be usefully approached from two complementary points of view: via the anomaly-inflow terms for background fields in the bulk, or in terms of the anomalous boundary degrees of freedom. In this paper we mostly take the former, bulk point of view.
@@
For our purposes in this paper, the most important thing about the Chern-Simons SPT in (2.2) is the quantization of its coefficient $k\in{\mathbb{Z}}$. If we imagine dialing the continuous coupling constants of the theory while preserving the $U(1)$ symmetry, and without closing the bulk gap,555555 Recall that a boundary, if present, necessarily hosts gapless degrees of freedom to cancel the bulk anomaly. then $k$ does not change. There are only two ways in which $k$ can change:
@@ 2.2.1 Basic Properties
  • The $\theta_{f}=\pi$ SPT represents the anomaly-inflow action for the parity anomaly [109, 110] in 2+1 dimensions: on a manifold $M_{4}$ with boundary $M_{3}=\partial M_{4}$, the $\theta_{f}=\pi$ SPT can be integrated by parts to produce a Chern-Simons term (2.2) with level $k={1\over 2}$ on $M_{3}$.565656 This in turn gives rise to a half-integer Hall conductance in units of $e^{2}/h$, as discussed in section 2.2.3 below. Since this term flips sign under either ${\mathsf{T}}$ or $\sf C{\mathsf{T}}$, the partition function of the boundary degrees of freedom on $M_{3}$ must compensate by transforming as follows,
@@ 2.2.2 Example: Massive Dirac Fermion
Here $\alpha=1,2$ is a left-handed Weyl spinor index.575757 We use the conventions of Wess and Bagger [65] for 2-component Weyl spinors. See also [74]. Hermitian conjugation exchanges left-handed undotted Weyl spinors with dotted right-handed Weyl spinors. Thus the hermitian conjugate fields are
@@
    Thanks to the standard axial-vector triangle anomaly, the $U(1)_{\text{axial}}$ symmetry is anomalous if we turn on a background field $A$ for the vector-like $U(1)$ flavor symmetry.595959 For now we ignore gravitational backgrounds; we will include them below. This leads to the following transformation rule for the partition function,606060 This formula can be derived using the standard descent formalism, which relates triangle anomalies to index theorems; see for instance section 2 of [11] for a recent pedagogical discussion with references.
@@ 2.2.3 Detecting the $\theta_{f}=\pi$ SPT via 2+1d Quantum Hall Conductance
Consider the codimension-one defect that generates the ${\mathsf{T}}$-symmetry, oriented along a 2+1d submanifold of spacetime. Since this defect implements the action of ${\mathsf{T}}$, it maps the $\theta_{f}=\pi$ SPT on one side to $\theta_{f}=-\pi$, leading to an overall jump by $\Delta\theta_{f}=2\pi$ across the wall. This has the effect of decorating the ${\mathsf{T}}$-defect by a $U(1)$ Chern-Simons term with level $k=1$,616161 See [120, 52, 121, 122, 123, 124] for a discussion of decorated domain walls. which in turn implies a non-zero quantum Hall conductance $\sigma_{xy}=1$, in units of $e^{2}/h$ (with $e$ the electron charge, and $h$ Planck’s constant), see for instance [39, 125].
@@ 2.3.1 Basic Properties
On a manifold with boundary this SPT can be integrated by parts to produce what is effectively a level-${1\over 2}$ gravitational Chern-Simons term,626262 This in turn gives rise to a thermal Hall conductance $\kappa_{xy}=1/4+n/2$ with $n\in{\mathbb{Z}}$, in units of $\pi^{2}k_{B}^{2}T/3h$, where $T$ is temperature, $h$ is Planck’s constant, and $k_{B}$ is Boltzmann’s constant. This is further discussed in section 2.3.3 below. whose flip under ${\mathsf{T}}$ or ${\mathsf{C}}{\mathsf{T}}$ transforms the partition function by a properly quantized gravitational Chern-Simons counterterm. This is the purely gravitational version of the parity anomaly.
@@ 2.3.2 Example: Massive Weyl Fermion
Note that a single Weyl fermion does not have charge-conjugation symmetry.636363 The ${{\mathsf{C}}{\mathsf{P}}{\mathsf{T}}}$ theorem also guarantees invariance under ${\mathsf{C}}{\mathsf{P}}:\psi_{\alpha}(t,\vec{x})\to i\overline{\psi}^{\dot{\alpha}}(t,-\vec{x})$. As before, we choose counterterms so that $m>0$ corresponds to the trivial SPT.
@@ 2.5.2 Partition Functions and Gapless SPTs
One way to diagnose gapped SPT phases is via their partition function in the presence of background fields on euclidean spacetime manifolds. Here we will make some related observations in the context of gapless SPTs.646464 We are grateful to Ryan Thorngren for discussions about the content of this subsection, and for explaining some of the intuitive arguments underlying [87]. We distinguish three cases:
@@ 2.5.3 Gapless SPT Phase with a $U(1)$ Nambu-Goldstone Boson
In this paper we will encounter gapless phases that contain a NGB for a spontaneously broken flavor symmetry $U(1)_{B}$ with background field $B$. In addition, there can be SPTs for a distinct, unbroken $U(1)$ flavor symmetry with background field $A$, as well as gravity, both protected by an unbroken time-reversal symmetry. Let us take the $U(1)_{B}$ Goldstone boson $\phi$ to be a dimensionless scalar of periodicity $\phi\sim\phi+2\pi$. The shift symmetry of $\phi$, with current $f_{\phi}d\phi$ coupling to $B$, is the broken $U(1)_{B}$ symmetry.656565 Here $f_{\phi}$ is the decay constant of $\phi$. In addition, $\phi$ has an unbroken, generally emergent, $U(1)_{C}^{(2)}$ two-form winding symmetry, whose current $\sim*d\phi$ couples to a three-form gauge field $C^{(3)}$ with field strength $H^{(4)}=dC^{(3)}$. All together the low-energy Lagrangian for $\phi$ is,
@@
Note that the broken $U(1)_{B}$ symmetry and the unbroken $U(1)_{C}^{(2)}$ winding symmetry have a mixed ’t Hooft anomaly. In the presentation (2.53), the theory is manifestly invariant under $C^{(3)}$ background gauge transformations,666666 If we add to the action a counterterm $\sim\int_{M_{4}}B\wedge C^{(3)}$ then a $C^{(3)}$ gauge transformation leads to a shift $\sim dB$. Clearly such a shift cannot eat an SPT involving only the $U(1)$ background field $A$ or gravity. but a $U(1)_{B}$ background gauge transformation $B\to B+d\lambda_{B}$ shifts
@@
Note that both ${1\over 2\pi}H^{(4)}$ and ${1\over 8\pi^{2}}F\wedge F$ have integral periods. However, substituting back into (2.54) we find a non-vanishing mixed anomaly between the broken $U(1)_{B}$ symmetry and the unbroken $U(1)$, in accord with the general discussion in section 2.5.1.676767 In this example, the mixed ’t Hooft anomaly between the $U(1)$ flavor symmetry and the broken $U(1)_{B}$ symmetry entirely comes from the mixed ’t Hooft anomaly between the $U(1)_{C}^{(2)}$ two-form symmetry and the $U(1)_{B}$ symmetry. Completely analogously, comparing with (2.27), we find that we can absorb $\theta_{g}=\pi$ SPT by choosing
@@ 3.1 Fields and Lagrangian
Here $\alpha=1,2$ is a left-handed Weyl spinor index,686868 We use the conventions of Wess and Bagger [65] for 2-component Weyl spinors. See [74] for a presentation of QCD in these conventions. which we will generally suppress. The raised and lowered $a$-indices are fundamental and anti-fundamental $SU(N)_{c}$ color indices, respectively, so that the pair $(\psi^{a}_{i},\chi_{a}^{i})$ transforms as $({\bf N},\overline{\bf N})$ of the $SU(N)_{c}$ gauge group. Likewise, the lowered and raised $i$-indices are fundamental and anti-fundamental $SU(N_{f})$ flavor indices, and $(\psi^{a}_{i},\chi_{a}^{i})$ transforms as $({\bf N_{f}},\overline{\bf N_{f}})$ under $SU(N_{f})$.
@@
In this paper we will mostly discuss the case of $N_{f}$ degenerate quarks, with a common real Dirac mass $m_{q}\in{\mathbb{R}}$, and vanishing theta-angle, $\theta_{c}=0$, for the dynamical $SU(N)_{c}$ gauge field.696969 An anomalous $U(1)_{\text{axial}}$ transformation can be used to eliminate $\theta_{c}$ in favor of a (generally) complex mass $m_{q}\in{\mathbb{C}}$. Below we will additionally impose time-reversal symmetry, which requires real $m_{q}\in{\mathbb{R}}$. The Lagrangian in Lorentzian signature is then given by
@@ 3.2 Continuous Symmetries
The continuous gauge and global symmetries that act faithfully on the quark fields are707070 See for instance section 4.1 of [114] for a related discussion.
@@
Here $SU(N)_{c}$ is the gauge group and $U(N_{f})_{f}$ is the vector-like flavor symmetry preserved by the quark mass.717171 Whenever we wish to highlight that something pertains to color or flavor we will use the subscripts $c$ and $f$ respectively, though we will omit them if confusion is unlikely to arise. Note that
@@
    Such a theory is bosonic: even though it is formulated using fermionic quark fields, all gauge invariant local operators are bosons and the theory can be studied on non-spin manifolds.727272 This utilizes the symmetry group $(SU(N)_{c}\times\text{Spin}(4)_{\text{Lorentz}})/{\mathbb{Z}}_{2}^{F}$ that arises when $N$ is even. On oriented Riemannian manifolds $M_{4}$ without a spin structure, the second Stiefel-Whitney class $w_{2}(M_{4})\in H^{2}(M_{4},{\mathbb{Z}}_{2})$ of the tangent bundle is non-trivial, while the gauge fields are connections on an $SU(N)_{c}/{\mathbb{Z}}_{2}$ bundle. The ${\mathbb{Z}}_{2}^{F}$ quotient in the symmetry group implies that the class that measures the obstruction to lifting the $SU(N)_{c}/{\mathbb{Z}}_{2}$ gauge bundle to an $SU(N)_{c}$ bundle is precisely given by $w_{2}(M_{4})$.
@@
    In particular this implies that $(-1)^{F}$ is identified with $(-1)\in U(1)_{B}\subset G_{f}$, so that fundamental baryons and anti-baryons with $B=\pm 1$ are necessarily fermions. This is consistent with the fact that they contain an odd number $N$ of constituent quarks. This identification further implies that the theory can be placed on non-spin manifolds, as long as the $U(1)_{B}$ background gauge field is taken to be a spinc connection.737373 In general this leads to additional information not accessible on spin manifolds, see e.g. [111, 140, 141, 142, 143].